Calculate the following binomial probability by either using one of the binomial probability tables, software, or a calculator using the formula below. Round your answer to 3 decimal places.

P(x | n, p) =

n!
(nx)! x!

· px · qnx    where    q = 1 − p

P(x > 10, n = 15, p = 0.8)

=
Answer #1

We have given here,

n=15 and p=0.8

P(x>10)

=P(x=11) + P(X=12) +P(x=13) +P(x=14) +P(x=15)

=0.188+0.25+0.231+0.132+0.035

=0.836

x P(x)
0 0
1 0
2 0
3 0
4 0
5 0
6 0.001
7 0.003
8 0.014
9 0.043
10 0.103
11 0.188
12 0.25
13 0.231
14 0.132
15 0.035

Excel command for exact probability BINOMDIST(x,n,p,FALSE)

For example if x=12

=BINOMDIST(12,15,0.8,FALSE)